A hot-air balloon of mass 500 kg is descending at a speed of 5 ms-1 with an acceleration of 0.2 ms-2.

(i) What mass of ballast must be thrown out to reduce the acceleration to zero?

(ii) If twice this amount of ballast is thrown out, how much further will the balloon descend before it starts to climb?

Answers: (i) 10 kg (ii) 60 m
Answer #1

o ifm is the mass y balloon and it is descending down with acceleration f. then From 2nd law Newtas the equation 1 to tif 7 gm- 510 x 0:2 16 tokg m=lokg From ean (ii) Fo = (m-mig FR = 1500-10) X10 = 4900 N. med at (M-2mJg whene am mass is thrmon out